# This is a preprint submitted to: Archives of Computational Methods in Engineering

### v1

@@ 6.2.3 Autoencoders
Autoencoders (AEs) are a type of artificial neural networks that are used primarily for the task of data compression and decompression through a dual-component architecture [91] (see Fig. 18). This architecture consists of two main functions: (1) an encoder function $\mathcal{E}$ maps the data $\mathbf{d}$ to compress data ${\mathbf{x}}$; (2) a decoder function $\mathcal{D}$ maps from the compressed data ${\mathbf{x}}$ back to $\hat{\mathbf{d}}$. This operation is performed by setting the same number of neurons $3L$ in the input and output layer and constraining the hidden layer to have $N<M$ neurons.
@@
DAEs have been effectively used for marine engineering problems of single- [88] and multi-hulls [95], as well as for aerodynamic optimization of RAE 2822 foil [96].
Despite their versatility, AEs require careful tuning of parameters, including the number of neurons $N$ in the hidden layer, which corresponds to the intrinsic dimensionality of the data. The choice of this dimensionality, which is not known a priori, is crucial for balancing detail in the reconstructed output against the simplicity of the model. Furthermore, the effectiveness of an AE can be significantly influenced by the initial weights and the configuration of the network layers. Continued advancements in AE technology, especially with the integration of regularization techniques and advanced training algorithms, promise to enhance their applicability in more complex and computationally demanding shape optimization tasks.
@@ 6.2.4 Generative topographic mapping
This model not only helps in reducing the dimensionality but also in understanding the underlying structure of the shape space. By optimizing the parameters $\mathbf{W}$, GTM learns a manifold that best represents the distribution of the actual data points, thereby facilitating a smoother interpolation and visualization of the shape parameters’ effects. The fitting process involves maximizing the likelihood of the data under the model, typically by employing an expectation-maximization algorithm, which adjusts $\mathbf{W}$ and other model parameters to better fit the observed data.
@@
In conclusion, while GTM presents a powerful tool for shape optimization by enabling effective dimensionality reduction and intuitive visualization of complex relationships, it requires careful management and understanding of its limitations. It has been applied for external [97] and internal [98] aerodynamics problems, nevertheless, its success in practical applications hinges on the balance between model complexity and computational feasibility, making it essential for users to thoroughly evaluate the model’s behavior in the context of specific optimization tasks.
@@ 7 Discussion and Conclusions
Geometry-based PCA and its variants are found to be the most used, even for the implementation simplicity, nevertheless they can not be effective since no physical insight is taken into consideration in the reduced design space representation. To overcome this issue, further advancement has been achieved by integrating physical information within the design-space preliminary assessment. Physics-informed methods represent a promising frontier, incorporating domain-specific knowledge to guide the dimensionality reduction process. By integrating physical principles, these methods aim to ensure that the reduced design spaces remain meaningful and relevant to the optimization objectives.
In conclusion, this survey illustrates the dynamic and multifaceted landscape of design-space dimensionality reduction techniques in shape optimization. Future research directions include the further development of physics-informed approaches, the exploration of novel nonlinear dimensionality reduction algorithms [99], as well as the use of more sophisticated space reduction [100], and the integration of these methods into comprehensive optimization frameworks. The continued evolution of these techniques will undoubtedly enhance their applicability and effectiveness in shape optimization and beyond, offering new opportunities for innovation in multidisciplinary optimization challenges, including also structural [101, 102] and topology [103, 104] optimization.
@@ Acknowledgments
The work was conducted in collaboration with the NATO task group AVT-404 on “Enhanced Design Processes of Military Vehicles through Machine Learning Methods”. The authors are grateful to the US Office of Naval Research Global for its support through grants N62909-11-1-7011 and N62909-21-1-2042, and acknowledge the support of the Italian Ministry of University and Research (MUR) through the PRIN 2022 BIODRONES project, code 20227JNM52 (CUP: 853023005560006).

### v2

@@ 6.2.3 Autoencoders
Autoencoders (AEs) are a type of artificial neural networks that are used primarily for the task of data compression and decompression through a dual-component architecture [127] (see Fig. 25). This architecture consists of two main functions: (1) an encoder function $\mathcal{E}$ maps the data $\mathbf{d}$ to compress data ${\mathbf{x}}$; (2) a decoder function $\mathcal{D}$ maps from the compressed data ${\mathbf{x}}$ back to $\hat{\mathbf{d}}$. This operation is performed by setting the same number of neurons in the input and output layer and constraining the hidden layer to have $N<M$ neurons.
@@
DAEs have been effectively employed for marine engineering problems involving single-hull [123, 131] and multi-hull designs [132]. In aerodynamic optimization contexts, DAEs have been successfully applied to the design of RAE 2822 airfoils [133], transonic wing optimization [134], aeroacoustic shape optimization of wind turbine blades [135], and rotorcraft airfoil aeroacoustic optimization tasks [88]. Recent studies have further advanced the field by integrating convolutional autoencoders for improved ship hull shape transformations [131] and combining autoencoder architectures with physics-informed neural networks for efficient aerodynamic airfoil optimization [136], demonstrating their versatility and robustness across diverse optimization scenarios. Additionally, graph variational autoencoders have been recently employed for direct 3D aerodynamic shape optimization, leveraging their ability to efficiently represent complex, non-parametric 3D geometries and providing compact and smooth latent spaces suitable for advanced exploration via evolutionary algorithms [137].
Figure 27: Block diagram for the solution of shape optimization problems through GTM
Despite their versatility, AEs require careful tuning of parameters, including the number of neurons $N$ in the hidden layer, which corresponds to the intrinsic dimensionality of the data. The choice of this dimensionality, which is not known a priori, is crucial for balancing detail in the reconstructed output against the model’s simplicity. Furthermore, the effectiveness of an AE can be significantly influenced by the initial weights, the configuration of the network layers, and the training phase (see Fig. 26). Continued advancements in AE technology, especially with the integration of regularization techniques and advanced training algorithms, promise to enhance their applicability in more complex and computationally demanding shape optimization tasks.
@@ 6.2.4 Generative topographic mapping
This model not only helps in reducing the dimensionality but also in understanding the underlying structure of the shape space. By optimizing the parameters $\mathbf{W}$, GTM learns a manifold that best represents the distribution of the actual data points, thereby facilitating a smoother interpolation and visualization of the shape parameters’ effects. The fitting process involves maximizing the likelihood of the data under the model, typically by employing an expectation-maximization algorithm, which adjusts $\mathbf{W}$ and other model parameters to better fit the observed data (see Fig. 27).
@@
It has been applied for external [138, 139] and internal [140] aerodynamics problems, nevertheless, its success in practical applications hinges on the balance between model complexity and computational feasibility, making it essential for users to thoroughly evaluate the model’s behavior in the context of specific optimization tasks. Recently [141], an optimized GTM has been proposed to overcome some of these issues, introducing negative log-likelihood of data to measure the fitting quality between the nonlinear manifold with different dimensions and the high-dimensional space, as well as a variable-fidelity sample filtration to generate an efficient training set.
In conclusion, while GTM presents a powerful tool for shape optimization by enabling effective dimensionality reduction and intuitive visualization of complex relationships, it requires careful management and understanding of its limitations.
@@ 7 Discussion and Conclusions
Geometry-based PCA and its variants are found to be the most used, even for the implementation simplicity, nevertheless they can not be effective since no physical insight is taken into consideration in the reduced design space representation. To overcome this issue, further advancement has been achieved by integrating physical information within the design-space preliminary assessment. Physics-informed methods represent a promising frontier, incorporating domain-specific knowledge to guide the dimensionality reduction process. By integrating physical principles, these methods aim to ensure that the reduced design spaces remain meaningful and relevant to the optimization objectives [142].
In conclusion, this survey illustrates the dynamic and multifaceted landscape of design-space dimensionality reduction techniques in shape optimization. Future research directions include the further development of physics-informed approaches, the exploration of novel nonlinear dimensionality reduction algorithms [143], as well as the use of more sophisticated space reduction [144], and the integration of these methods into comprehensive optimization frameworks. Promising extensions of traditional PCA to nonlinear settings, such as principal geodesic analysis, have recently emerged, employing advanced manifold learning methods like Grassmannian manifolds and separable shape tensors to capture complex geometric variations that linear PCA cannot adequately represent [145, 146]. The continued evolution of these techniques will undoubtedly enhance their applicability and effectiveness in shape optimization and beyond, offering new opportunities for innovation in multidisciplinary optimization challenges, including also structural [147, 148] and topology [149, 150] optimization.
@@ Acknowledgments
The work was conducted in collaboration with the NATO-AVT-404 Research Task Group on “Enhanced Design Processes of Military Vehicles through Machine Learning Methods”. The authors are grateful to the US Office of Naval Research Global for its support through grants N62909-11-1-7011 and N62909-21-1-2042, and acknowledge the support of the Italian Ministry of University and Research (MUR) through the PRIN 2022 BIODRONES project, code 20227JNM52 (CUP: 853023005560006).